# 25mm25mm25mm26mm

### v2

@@ 3.2. Local laws for multiple resolvents
The main technical result of this manuscript are local laws for the product of two Green functions $G^{z_{1}}(\mathrm{i}\eta_{1})AG^{z_{2}}(\mathrm{i}\eta_{2})$, with different parameters $z_{i}\in{\mathbb{C}}$ and $\eta_{i}\in{\mathbb{C}}~{}(i=1,2)$, which yield the overlap bounds of the eigenvectors stated in Theorem 2.2. Before we state the local laws, we first introduce the deterministic approximation matrices of $G^{z_{1}}(\mathrm{i}\eta_{1})AG^{z_{2}}(\mathrm{i}\eta_{2})$, i.e.,
@@
$\mathcal{B}_{12}:=\mathcal{B}_{12}(z_{1},\omega_{1},z_{2},\omega_{2})=1-M^{z_{1}}(\omega_{1})\mathcal{S}[\cdot]M^{z_{2}}(\omega_{2}).$ (3.17)
We further define the deterministic matrix of $\mathrm{Im}\,G^{z_{1}}(\mathrm{i}\eta_{1})A_{1}\mathrm{Im}\,G^{z_{2}}(\mathrm{i}\eta_{2})$ as a fourfold linear combination of $M^{A}_{12}(\pm\eta_{1},z_{1},\pm\eta_{2},z_{2})$, using the identity $\mathrm{Im}\,G=\frac{1}{2\mathrm{i}}(G-G^{*})$ twice, denoted by
$\widehat{M}_{12}^{A_{1}}=\widehat{M}_{12}^{A_{1}}(\mathrm{i}\eta_{1},z_{1},\mathrm{i}\eta_{2},z_{2}).$ (3.18)
@@
uniformly for any $|z_{i}|<10$ and $|\eta_{i}|<10$, with $\rho^{*}:=\max_{i=1,2}|\rho_{i}|$, $\eta_{*}:=\min_{i=1,2}|\eta_{i}|$ and the following parameters:
@@
$\widehat{\gamma}=\widehat{\gamma}(\mathrm{i}\eta_{1},z_{1},\mathrm{i}\eta_{2},z_{2}):=\widehat{\gamma}:=|z_{1}-z_{2}|^{2}+\rho_{1}|\eta_{1}|+\rho_{2}|\eta_{2}|+\left(\frac{\eta_{1}}{\rho_{1}}\right)^{2}+\left(\frac{\eta_{2}}{\rho_{2}}\right)^{2}.$ (3.24)
@@
Fix small constants $\epsilon,\tau>0$. Then, for any deterministic matrices $\|A_{i}\|\lesssim 1$ and vectors $\|{\bm{x}}\|,\|{\bm{y}}\|\lesssim 1$, the following local laws
@@ 3.3. Improved local law at the edge
Fix small constants $\epsilon,\tau>0$ with $100<\tau<\epsilon/$ 10. Then, for any deterministic matrices $\|A_{i}\|\lesssim 1$, the following local law
@@
Fix a small constant $\tau>0$. Then there exists a small $100<\omega_{c}<\tau/$ 10 such that, for any $z_{i}\in{\mathbb{C}}~{}(i=1,2)$ with $|z_{i}|\leq 1+n^{-1/2+\tau}$, the following holds with very high probability,
@@
We conclude this section by showing that the bound (3.27) readily implies those in Theorem 2.2 and Corollary 2.4.
@@
Let $\eta_{\mathrm{f}}=\eta_{\mathrm{f}}(z)$ denote the local fluctuation scale of the eigenvalues of $H^{z}$ close to zero, which, for $|z|\leq 1+n^{-1/2+\tau}$ with a small fixed $\tau>0$, is implicitly defined by
@@
Note that $\eta_{\mathrm{f}}\sim n^{-1}$ when $|z|<1$ and $\eta_{\mathrm{f}}\sim n^{-3/4}$ when $|z|=1$. Fix a small $\xi>0$, and define $\eta_{i}:=n^{\xi}\eta_{\mathrm{f}}(z_{i})$, for $z_{i}\in{\mathbb{C}}$. Then, by (3.27) and (3.21), we obtain
@@
with very high probability. Here $\rho_{i}:=\pi^{-1}|\mathrm{Im}\,m^{z_{i}}(\mathrm{i}\eta_{i})|$ and $\widehat{\gamma}=\widehat{\gamma}(\mathrm{i}\eta_{1},z_{1},\mathrm{i}\eta_{2},z_{2})$. We point out that to obtain the supremum bound in (3.29) we used (3.27) and a standard grid argument using the Lipschitz continuity of $z\mapsto\mathrm{Im}\,G^{z}$.
@@
with very high probability, for some small constant $100<\omega_{c}<\xi/$ 10.
@@
$\langle\mathrm{Im}\,G^{z_{1}}(\mathrm{i}\eta_{1})\mathrm{Im}\,G^{z_{2}}(\mathrm{i}\eta_{2})\rangle=\frac{4}{n}\sum_{j,k=1}^{n}\frac{\eta_{1}\eta_{2}\big{(}|\langle\mathbf{u}^{z_{1}}_{j},\mathbf{u}_{k}^{z_{2}}\rangle|^{2}+|\langle\mathbf{v}^{z_{1}}_{j},\mathbf{v}_{k}^{z_{2}}\rangle|^{2}\big{)}}{\big{(}(\lambda^{z_{1}}_{j})^{2}+\eta_{1}^{2}\big{)}((\lambda^{z_{2}}_{k})^{2}+\eta_{2}^{2}\big{)}}.$ (3.31)
Then, on the very high probability event on which (3.29)–(3.30) hold, we obtain (recall $\eta_{i}:=n^{\xi}\eta_{\mathrm{f}}(z_{i})$)
$\begin{split}&\sup_{|z_{1}|,|z_{2}|\leq 1+n^{-1/2+\tau}}(n|z_{1}-z_{2}|^{2}+1)\big{[}|\langle\mathbf{u}^{z_{1}}_{1},\mathbf{u}_{1}^{z_{2}}\rangle|^{2}+|\langle\mathbf{v}^{z_{1}}_{1},\mathbf{v}_{1}^{z_{2}}\rangle|^{2}\big{]}\\ &\qquad\leq\sup_{|z_{1}|,|z_{2}|\leq 1+n^{-1/2+\tau}}(n|z_{1}-z_{2}|^{2}+1)(n\eta_{1}\eta_{2})\langle\mathrm{Im}\,G^{z_{1}}(\mathrm{i}\eta_{1})\mathrm{Im}\,G^{z_{2}}(\mathrm{i}\eta_{2})\rangle\\ &\qquad\lesssim\sup_{|z_{1}|,|z_{2}|\leq 1+n^{-1/2+\tau}}\frac{|z_{1}-z_{2}|^{2}+n^{-1}\normalcolor}{\widehat{\gamma}}(n\eta_{1}\rho_{1})(n\eta_{2}\rho_{2})\lesssim n^{2\xi},\end{split}$ (3.32)
with very high probability, where in the last inequality we used (3.24). We conclude the proof by noticing that, by the basic relation (1.6), (3.32) immediately implies (2.4)–(2.3) by possibly redefining $\xi$ and choosing $z_{1}=\sigma_{i}$ and $z_{2}=\sigma_{j}$ and noticing that with very high probability there are no eigenvalues outside $|z|\leq 1+n^{-1/2+\tau}$ (see [9, Theorem 2.1]). ∎
@@
Assume that $z,z\textquoteright\in{\mathbb{C}}$ are in the spectrum of $X$. Without loss of generality, we may write $\sigma_{i}=z$ and $\sigma_{j}=z\textquoteright$ ’, i.e., $X\mathbf{r}_{i}=z\mathbf{r}_{i}$, $\mathbf{l}^{*}_{i}X=z\mathbf{l}^{*}_{i}$ and similarly for the index $j$ and $z\textquoteright$ ’. By the definition of $H^{z}$ in (3.1) and the fundamental relation (1.6), zero is an eigenvalue of $H^{z}$ and $(\mathbf{r}^{*}_{i},\mathbf{l}^{*}_{i})^{*}\in{\mathbb{C}}^{2n}$ lies in the kernel of $H^{z}$, similarly for $j$, $z\textquoteright$ ’. Thus conditioning on that $\sigma_{i}=z$ and $\sigma_{j}=z\textquoteright$ ’, we reduce the eigenvector overlap $\mathcal{O}_{ij}$ of $X$ to the overlaps of the eigenvectors of the Hermitised matrix. Then, combining the very high probability bound from (3.32) with various bounds on the expectation of the diagonal overlaps from [44, Theorems 2.7, 2.9], by a simple Cauchy-Schwarz inequality, we obtain the desired results in Corollary 2.4. The additional conditions stem from the requirements in [44]. ∎
@@ 4. Zig flow
In this section we show that if certain local laws (see [27, Theorem 5.2]) hold for $\eta\sim 1$, then they can be propagated down to $\eta\gg n^{-1}$ using the Ornstein–Uhlenbeck and characteristic flows. We point out that some parts of this section, like the definition and properties of the characteristics flow and the analysis of the flow in the averaged case, are very close to [31, Section 5]; however, we decided to present the details here for the convenience of the reader. Consider the Ornstein–Uhlenbeck flow
@@
Throughout this section we consider spectral parameters being $2n\times 2n$ block–constant matrices, i.e. we consider $\Lambda$’s consisting of four $n\times n$ blocks each one of them being a constant multiple of the $n$–dimensional identity matrix. The main object of interest within this section is the resolvent $(W_{t}-\Lambda_{t})^{-1}$, with $W_{t}$ from (4.2) and $\Lambda_{t}$ being the solution of the characteristic flow:
@@
Fix any $T\leq 1$ and pick $\Lambda$ of the form (4.3) with $|\eta|>0$ and $\lVert\Lambda\rVert\leq 1$. Then there exist $\Lambda_{0}$ with $\lVert\Lambda_{0}\rVert\lesssim 1$ and $T^{*}(\eta_{0},z_{0})>T$, and the solution $\Lambda_{t}$ of (4.4) with initial condition $\Lambda_{0}$ is so that $\Lambda_{T}=\Lambda$. Furthermore, there exists an $N$–independent constant $c_{*}>0$ such that $\mathrm{dist}(\mathrm{i}\eta_{0},\mathrm{supp}(\rho^{z_{0}}))\geq c_{*}T$.
@@
Fix small $n$–independent constants $\tau,\omega_{*},\epsilon>0$, and for $i=1,2$ fix spectral parameters
@@
with $|z_{i,0}|\leq 10$, $|\eta_{i,0}|\leq\omega_{*}$. For $t\leq T^{*}$ let $\Lambda_{i,t}$ be the solution of (4.4) with initial condition $\Lambda_{i,0}$. Let $G_{i,t}:=(W_{t}-\Lambda_{i,t})^{-1}$, and let $M_{12,t}^{A}$ be defined in (4.10). Define the following control parameters and their time–dependent versions:
@@
Assume that for some small $100<\xi\leq\epsilon/$ 10, with very high probability, it holds
@@
uniformly in $\lVert A_{i}\rVert\leq 1$ and for any $|\eta_{i}|\in[|\eta_{i,0}|,n^{100}]$, where $\ell_{0},\gamma_{0}$ are evaluated at $(\eta_{1},z_{1,0},\eta_{2},z_{2,0})$. Then,
@@
Fix small $n$–independent constants $\tau,\omega_{*},\epsilon>0$, and for $i=1,2$ fix spectral parameters $\Lambda_{i,0}$ as in (4.12). Fix an exponent $b\in[0,1]$. For $t\leq T^{*}$ let $\Lambda_{i,t}$ be the solution of (4.4) with initial condition $\Lambda_{i,0}$. Let $G_{i,t}:=(W_{t}-\Lambda_{i,t})^{-1}$, and let $M_{12,t}^{A}$, $M_{121,t}^{A_{1},A_{2}}$ be defined in (4.10) and (4.11), respectively. Let $\ell_{t},\gamma_{t}$ be from (4.13), and assume that for some small $100<\xi\leq\epsilon/$ 10, with very high probability, it holds
@@
Note that the estimate on the three-resolvent isotropic chain is far from optimal, it shows only that the random chain can be bounded by the size of its deterministic approximation $M_{121}$, see (3.22). This will, however, be sufficient for our purpose since for the quadratic variation we anyway need only a bound on the random chain itself. We are now ready to present the proofs of Propositions 4.34.4. We first present the proof in the averaged case and then consider the isotropic case.
@@
Here $\partial_{ab}$ denotes the directional derivative in the direction $(W_{t})_{ab}$ and $\sum_{i\neq j}$ denotes a summation over the pairs $(i,j)\in\{(1,2),(2,1)\}$. We point out that here we also used the following lemma describing the evolution of the deterministic approximation for the product of two resolvents:

### v3

@@ 3.2. Local laws for multiple resolvents
The main technical result of this manuscript are local laws for the product of two Green functions
$G^{z_{1}}(\mathrm{i}\eta_{1})AG^{z_{2}}(\mathrm{i}\eta_{2}),\qquad A\in{\mathbb{C}}^{2n\times 2n},\quad G^{z_{i}}(\mathrm{i}\eta_{i})\in{\mathbb{C}}^{2n\times 2n},~i=1,2,$
with different parameters $z_{i}\in{\mathbb{C}}$ and $\eta_{i}\in{\mathbb{R}}\setminus\{0\}~(i=1,2)$, which yield the overlap bounds of the eigenvectors stated in Theorem 2.2. Before we state the local laws, we first introduce the deterministic approximation matrices of $G^{z_{1}}(\mathrm{i}\eta_{1})AG^{z_{2}}(\mathrm{i}\eta_{2})$, i.e.,
@@
$\mathcal{B}_{12}:=\mathcal{B}_{12}(z_{1},\omega_{1},z_{2},\omega_{2})=1-M^{z_{1}}(\omega_{1})\mathcal{S}[\cdot]M^{z_{2}}(\omega_{2}),$ (3.17)
with the covariance operator $\mathcal{S}[\cdot]$ defined in (3.6); see also (4.10) later for the definition of $\mathcal{B}_{ij}$ with general $i,j\in{\mathbb{N}}$. Using the identity $\mathrm{Im}\,G=\frac{1}{2\mathrm{i}}(G-G^{*})$ twice and the definition of $M^{A}_{12}$ in (3.16), we further define the deterministic approximating matrix of $\mathrm{Im}\,G^{z_{1}}(\mathrm{i}\eta_{1})A\mathrm{Im}\,G^{z_{2}}(\mathrm{i}\eta_{2})$ as a fourfold linear combination of $M^{A}_{12}(\pm\mathrm{i}\eta_{1},z_{1},\pm\mathrm{i}\eta_{2},z_{2})$ as defined in (3.16), i.e.,
$\widehat{M}_{12}^{A}:=$ $\widehat{M}_{12}^{A}(\mathrm{i}\eta_{1},z_{1},\mathrm{i}\eta_{2},z_{2})$
$=$ $\frac{1}{4}\big(M^{A}_{12}(\mathrm{i}\eta_{1},z_{1},-\mathrm{i}\eta_{2},z_{2})+M^{A}_{12}(-\mathrm{i}\eta_{1},z_{1},\mathrm{i}\eta_{2},z_{2})-M^{A}_{12}(\mathrm{i}\eta_{1},z_{1},\mathrm{i}\eta_{2},z_{2})-M^{A}_{12}(-\mathrm{i}\eta_{1},z_{1},-\mathrm{i}\eta_{2},z_{2})\big).$ (3.18)
@@
uniformly for any $|z_{i}|<10$ and $0<|\eta_{i}|<10$, where $\eta_{*}:=\min_{i=1,2}|\eta_{i}|$ and $\rho^{*}:=\max_{i=1,2}|\rho_{i}|$ with $\rho_{i}:=\pi^{-1}|\mathrm{Im}\,m^{z_{i}}(\mathrm{i}\eta_{i})|$, and the following parameters:
@@
$\widehat{\gamma}=\widehat{\gamma}(\mathrm{i}\eta_{1},z_{1},\mathrm{i}\eta_{2},z_{2}):=|z_{1}-z_{2}|^{2}+\rho_{1}|\eta_{1}|+\rho_{2}|\eta_{2}|+\left(\frac{\eta_{1}}{\rho_{1}}\right)^{2}+\left(\frac{\eta_{2}}{\rho_{2}}\right)^{2}.$ (3.24)
@@
Fix small constants $\epsilon,\tau>0$. Then, for any deterministic matrices $\|A_{i}\|\lesssim 1$ and vectors $\|{\bm{x}}\|,\|{\bm{y}}\|\lesssim 1$, the following isotropic local law:
@@ 3.3. Improved local law at the edge
Fix small constants $\epsilon,\tau>0$ with $0<\tau<\epsilon/10$. Then, for any deterministic matrices $\|A_{i}\|\lesssim 1$, the following local law
@@
Fix a small constant $\tau>0$. Then there exists a small $0<\omega_{c}<\tau/10$ such that, for any $z_{i}\in{\mathbb{C}}~(i=1,2)$ with $|z_{i}|\leq 1+n^{-1/2+\tau}$, the following holds with very high probability,
@@
We conclude this section by showing that the bound (3.27) readily implies those in Theorem 2.2 and Corollary 2.5. Actually, Corollary 3.6 is obtained as a by-product in the proof of Theorem 2.2 (see (3.32) below).
@@
Let $\eta_{\mathrm{f}}=\eta_{\mathrm{f}}(z)$ denote the local fluctuation scale of the eigenvalues of $H^{z}$ close to zero, which, for $|z|\leq 1$ is implicitly defined by
@@
From (3.10), note that $\eta_{\mathrm{f}}\sim n^{-1}$ when $|z|<1$ and $\eta_{\mathrm{f}}\sim n^{-3/4}$ when $|z|=1$. For $|z|>1$ the density $\rho^{z}$ develops a gap $\Delta^{z}$ around zero, and one can see that in this case $\eta_{\mathrm{f}}\sim|\Delta^{z}|^{1/9}/n^{2/3}$ and $|\Delta^{z}|\sim(|z|^{2}-1)^{3/2}$ (see e.g. [30, Appendix A] and references therein). Fix a small $\xi>0$, and define $\eta_{i}:=n^{\xi}\eta_{\mathrm{f}}(z_{i})$, for $z_{i}\in{\mathbb{C}}$. Then, by (3.27) and (3.21), we obtain
@@
with very high probability. Here $\rho_{i}=\pi^{-1}|\mathrm{Im}\,m^{z_{i}}(\mathrm{i}\eta_{i})|$ and $\widehat{\gamma}=\widehat{\gamma}(\mathrm{i}\eta_{1},z_{1},\mathrm{i}\eta_{2},z_{2})$. We point out that to obtain the supremum bound in (3.29) we used (3.27) and a standard grid argument using the Lipschitz continuity of $z\mapsto\mathrm{Im}\,G^{z}$.
@@
with very high probability, for some small constant $0<\omega_{c}<\xi/10$. Note that for any $j\in[n]$, $\lambda^{z}_{j}$ is the $j$-th non-negative eigenvalue of $H^{z}$ in the non-deceasing order and $\lambda^{z}_{-j}=-\lambda^{z}_{j}$.
@@
$\langle\mathrm{Im}\,G^{z_{1}}(\mathrm{i}\eta_{1})\mathrm{Im}\,G^{z_{2}}(\mathrm{i}\eta_{2})\rangle=\frac{4}{n}\sum_{j,k=1}^{n}\frac{\eta_{1}\eta_{2}\big(|\langle\mathbf{u}^{z_{1}}_{j},\mathbf{u}_{k}^{z_{2}}\rangle|^{2}+|\langle\mathbf{v}^{z_{1}}_{j},\mathbf{v}_{k}^{z_{2}}\rangle|^{2}\big)}{\big((\lambda^{z_{1}}_{j})^{2}+\eta_{1}^{2}\big)((\lambda^{z_{2}}_{k})^{2}+\eta_{2}^{2}\big)},$ (3.31)
where $(\lambda^{z_{i}}_{j})_{j=1}^{n}$ are the non-decreasing singular values of $X-z_{i}$ with $i=1,2$, and $\{\mathbf{u}_{j}^{z_{i}}\}_{j=1}^{n}$, $\{\mathbf{v}_{j}^{z_{i}}\}_{j=1}^{n}$ are the corresponding left and right singular vectors normalized such that $\|\mathbf{u}_{j}^{z_{i}}\|^{2}=\|\mathbf{v}_{j}^{z_{i}}\|^{2}=1/2$. Then, on the very high probability event on which (3.29)–(3.30) hold, we obtain (recall $\eta_{i}:=n^{\xi}\eta_{\mathrm{f}}(z_{i})$)
$\begin{split}&\sup_{|z_{1}|,|z_{2}|\leq 1+n^{-1/2+\tau}}(n|z_{1}-z_{2}|^{2}+1)\big[|\langle\mathbf{u}^{z_{1}}_{1},\mathbf{u}_{1}^{z_{2}}\rangle|^{2}+|\langle\mathbf{v}^{z_{1}}_{1},\mathbf{v}_{1}^{z_{2}}\rangle|^{2}\big]\\ &\qquad\leq\sup_{|z_{1}|,|z_{2}|\leq 1+n^{-1/2+\tau}}(n|z_{1}-z_{2}|^{2}+1)(n\eta_{1}\eta_{2})\langle\mathrm{Im}\,G^{z_{1}}(\mathrm{i}\eta_{1})\mathrm{Im}\,G^{z_{2}}(\mathrm{i}\eta_{2})\rangle\\ &\qquad\lesssim\sup_{|z_{1}|,|z_{2}|\leq 1+n^{-1/2+\tau}}\frac{|z_{1}-z_{2}|^{2}+n^{-1}\normalcolor}{\widehat{\gamma}}(n\eta_{1}\rho_{1})(n\eta_{2}\rho_{2})\lesssim n^{4\xi},\end{split}$ (3.32)
with very high probability, where in the last inequality we used (3.24) to show that $(|z_{1}-z_{2}|^{2}+n^{-1})/\widehat{\gamma}\lesssim 1$ and we used that $n\eta_{i}\rho_{i}\lesssim n^{2\xi}$ by the definition of $\eta_{\mathrm{f}}$ above (3.29) (note that $\eta_{\mathrm{f}}$ has the property that $n\eta_{\mathrm{f}}\rho^{z}(\mathrm{i}\eta_{\mathrm{f}})\sim 1$).
Recall the basic relation (1.7) that $z\in\mathrm{Spec}(X)$ if and only if the least singular value $\lambda^{z}_{1}$ is zero. Moreover, if $z\in\mathrm{Spec}(X)$, then the left and right eigenvector of $X$ associated to the eigenvalue $z$ are exactly the left and right singluar vector of XzX-z. This implies that
$\sup_{i,j\in[n]}\big(n|\sigma_{i}-\sigma_{j}|^{2}+1\big)\left[\frac{\big|\langle\mathbf{r}_{i},\mathbf{r}_{j}\rangle\big|^{2}}{\lVert\mathbf{r}_{i}\rVert^{2}\lVert\mathbf{r}_{j}\rVert^{2}}+\frac{\big|\langle\mathbf{l}_{i},\mathbf{l}_{j}\rangle\big|^{2}}{\lVert\mathbf{l}_{i}\rVert^{2}\lVert\mathbf{l}_{j}\rVert^{2}}\right]$
$=$ $4\sup_{z_{1},z_{2}\in\mathrm{Spec}(X)}(n|z_{1}-z_{2}|^{2}+1)\big[|\langle\mathbf{u}^{z_{1}}_{1},\mathbf{u}_{1}^{z_{2}}\rangle|^{2}+|\langle\mathbf{v}^{z_{1}}_{1},\mathbf{v}_{1}^{z_{2}}\rangle|^{2}\big],$ (3.33)
where $\mathbf{u}^{z_{i}}_{1}$ and $\mathbf{v}^{z_{i}}_{1}~(i=1,2)$ are the left and right singular vectors associated to the least singular value of $X-z_{i}$, normalized such that $\|\mathbf{u}_{1}^{z_{i}}\|^{2}=\|\mathbf{v}_{1}^{z_{i}}\|^{2}=1/2$. Note that with very high probability the spectrum of $X$ lies inside the disk $|z|\leq 1+n^{-1/2+\tau}$ (see [9, Theorem 2.1]). Hence, with very high probability,
$\sup_{i,j\in[n]}\big(n|\sigma_{i}-\sigma_{j}|^{2}+1\big)\left[\frac{\big|\langle\mathbf{r}_{i},\mathbf{r}_{j}\rangle\big|^{2}}{\lVert\mathbf{r}_{i}\rVert^{2}\lVert\mathbf{r}_{j}\rVert^{2}}+\frac{\big|\langle\mathbf{l}_{i},\mathbf{l}_{j}\rangle\big|^{2}}{\lVert\mathbf{l}_{i}\rVert^{2}\lVert\mathbf{l}_{j}\rVert^{2}}\right]$
$\lesssim$ $\sup_{|z_{1}|,|z_{2}|\leq 1+n^{-1/2+\tau}}(n|z_{1}-z_{2}|^{2}+1)\big[|\langle\mathbf{u}^{z_{1}}_{1},\mathbf{u}_{1}^{z_{2}}\rangle|^{2}+|\langle\mathbf{v}^{z_{1}}_{1},\mathbf{v}_{1}^{z_{2}}\rangle|^{2}\big].$ (3.34)
This, together with (3.32), immediately implies (2.3)–(2.4) and concludes the proof. ∎
@@
Fix any small $\xi>0$. From (2.4), we know that for any $i,j\in[N]$, with very high probability,
$|\mathcal{O}_{ij}|\leq$ $\frac{n^{\xi}}{n|\sigma_{i}-\sigma_{j}|^{2}+1}\sqrt{\mathcal{O}_{ii}\mathcal{O}_{jj}}$
$\lesssim$ $\frac{n^{\xi}}{n|z_{1}-z_{2}|^{2}+1}\sqrt{\mathcal{O}_{ii}\mathcal{O}_{jj}},$ (3.35)
where the last step follows if we assume that the corresponding eigenvalues $\sigma_{i}\in\mathcal{D}^{\xi}_{z_{1}}$ and $\sigma_{j}\in\mathcal{D}^{\xi}_{z_{2}}$, with $|z_{1}-z_{2}|\geq n^{-1/2+\epsilon}$ and $\xi<\epsilon/10$. Moreover, from the local law mentioned in Remark 2.6, for any $z\in{\mathbb{C}}$, with very high probability,
$\#\{\sigma_{i}\in\mathrm{Spec}(X):\sigma_{i}\in\mathcal{D}^{\xi}_{z}\}\leq n^{3\xi}.$ (3.36)
We next recall various bounds on the expectation of the diagonal overlaps from [44, Theorems 2.7, 2.9]. For example, in the complex case, using [44, Eq (2.14)], there exists an event $\Xi$ with very high probability, i.e. ${\mathbf{P}}(\Xi)\geq 1-n^{-D}$, such that for any $z\in{\mathbb{C}}$,
${\mathbf{E}}\Big(\sum_{\sigma_{i}\in\mathcal{D}^{\xi}_{z}}|\mathcal{O}_{ii}|\cdot{\bf 1}_{\Xi}\Big)\lesssim n^{1+3\xi}.$ (3.37)
Combining the very high probability bound (3.3)-(3.36) with the expectation bound (3.37) and using a simple Cauchy-Schwarz inequality, we obtain the desired result in (2.7) for the complex case. The remaining statements of Corollary 2.5 follow similarly by combining (3.3)-(3.36) with the corresponding expectation bounds from Eq (2.15), Eq (2.17), and Eq (2.18) of [44], respectively. This finishes the proof. ∎
@@ 4. Zig flow
In this section we show that if certain local laws (see [27, Theorem 5.2]) hold for $\eta\sim 1$ (for simplicity in this introductory paragraph we focus on $|z|<1$), then they can be propagated down to $\eta\gg n^{-1}$ using the Ornstein–Uhlenbeck and characteristic flows. The main results of this section (Propositions 4.34.4) are the fundamental inputs to prove Theorem 3.4. In fact, they show that if the local laws from Theorem 3.4 hold on a global scale ($\eta\sim 1$), then they also hold down to the optimal $\eta$-scale ($\eta\gg 1/n$), at the price of adding a Gaussian component. This Gaussian component will then be removed in Section 5 using a Green’s Function Comparison Theorem (GFT). We point out that some parts of this section, like the definition and properties of the characteristics flow and the analysis of the flow in the averaged case, are very close to [31, Section 5]; however, we decided to present the details here for the convenience of the reader. Consider the Ornstein–Uhlenbeck flow
@@
Throughout this section we consider spectral parameters being $2n\times 2n$ block–constant matrices, i.e. we consider $\Lambda$’s consisting of four $n\times n$ blocks each one of them being a constant multiple of the nn–dimensional identity matrix. The main object of interest within this section is the resolvent $(W_{t}-\Lambda_{t})^{-1}$, with $W_{t}$ from (4.2) and $\Lambda_{t}$ being the solution of the characteristic flow:
@@
Fix any $T\leq 1$ and pick $\Lambda$ of the form (4.3) with $|\eta|>0$ and $\lVert\Lambda\rVert\leq 1$. Then there exist $\Lambda_{0}$ with $\lVert\Lambda_{0}\rVert\lesssim 1$ and $T^{*}(\eta_{0},z_{0})>T$, and the solution $\Lambda_{t}$ of (4.4) with initial condition $\Lambda_{0}$ is so that $\Lambda_{T}=\Lambda$. Furthermore, there exists an NN–independent constant $c_{*}>0$ such that $\mathrm{dist}(\mathrm{i}\eta_{0},\mathrm{supp}(\rho^{z_{0}}))\geq c_{*}T$.
@@
Fix small nn–independent constants $\tau,\omega_{*},\epsilon>0$, and for $i=1,2$ fix spectral parameters
@@
with $|z_{i,0}|\leq 10$, $|\eta_{i,0}|\leq\omega_{*}$. For $t\leq T^{*}$ let $\Lambda_{i,t}$ be the solution of (4.4) with initial condition $\Lambda_{i,0}$. Let $G_{i,t}:=(W_{t}-\Lambda_{i,t})^{-1}$, and let $M_{12,t}^{A}$ be defined in (4.10). Define the following control parameters and their time–dependent versions [^3] Recall that $\rho_{i}=\rho^{z_{i}}(\mathrm{i}\eta_{i})$, and so in the time dependent case we have $\rho_{i,t}=\rho^{z_{i,t}}(\mathrm{i}\eta_{i,t})$. For example, using these notations, $\ell_{t}:=\min_{i=1,2}\rho_{i,t}\eta_{i,t}$.:
@@
Assume that for some small $0<\xi\leq\epsilon/10$, with very high probability, with $G_{0}^{z}(\mathrm{i}\eta):=(H_{0}^{z}-\mathrm{i}\eta)^{-1}$, it holds
@@
uniformly in $\lVert A_{i}\rVert\leq 1$ and for any $|\eta_{i}|\in[|\eta_{i,0}|,n^{100}]$, where $\ell_{0},\gamma_{0}$ are evaluated at [^4] We point out that we require (4.14) for $(\eta_{1},z_{1,0},\eta_{2},z_{2,0})$, with $|\eta_{i}|\in[|\eta_{i,0}|,n^{100}]$, and not only for $\eta_{i}=\eta_{i,0}$. $(\eta_{1},z_{1,0},\eta_{2},z_{2,0})$. Then,
@@
Close to the edge of the spectrum, i.e. when $|z_{i}|\approx 1$, we instead have $\gamma\sim|z_{1}-z_{2}|+\eta_{1}^{2/3}+\eta_{2}^{2/3}$.
@@
Fix small nn–independent constants $\tau,\omega_{*},\epsilon>0$, and for $i=1,2$ fix spectral parameters $\Lambda_{i,0}$ as in (4.12). Fix an exponent $b\in[0,1]$. For $t\leq T^{*}$ let $\Lambda_{i,t}$ be the solution of (4.4) with initial condition $\Lambda_{i,0}$. Let $G_{i,t}:=(W_{t}-\Lambda_{i,t})^{-1}$, and let $M_{12,t}^{A}$, $M_{121,t}^{A_{1},A_{2}}$ be defined in (4.10) and (4.11), respectively. Let $\ell_{t},\gamma_{t}$ be from (4.13), and assume that for some small $0<\xi\leq\epsilon/10$, with very high probability, it holds
@@
Note that the estimate on the three-resolvent isotropic chain is far from optimal, it shows only that the random chain can be bounded by the size of its deterministic approximation $M_{121}$, see (3.22). This will, however, be sufficient for our purpose since for the quadratic variation we anyway need only a bound on the random chain itself. We are now ready to present the proofs of Propositions 4.34.4. We first present the proof in the averaged case and then consider the isotropic case. For the sake of brevity throughout this section (and in Section 4.1) we only focus on the complex case, all the additional terms appearing in the real case can be handled analogously, we thus omit the details.
@@
Here $\partial_{ab}$ denotes the directional derivative in the direction $(W_{t})_{ab}$, $B_{ab,t}$ denotes the (a,b)(a,b)–entry of the matrix Brownian motion $B_{t}$ from (4.1), and $\sum_{i\neq j}$ denotes a summation over the pairs $(i,j)\in\{(1,2),(2,1)\}$. We point out that here we also used the following lemma describing the evolution of the deterministic approximation for the product of two resolvents: